Acceleration due to gravity on the surface of venus

The mass of Venus is 81.5% that of the earth, and its radius is 94.9% that of the earth.
(a) Compute the acceleration due to gravity on the surface of Venus from these data. (Express your answer in terms of the acceleration due to gravity on theEarth.)
__ gE

(b) If a rock weighs 29.4 N on earth, what would it weigh at the surface of Venus?
___ N
Answer #1
The mass of Venus ( Mv = 81.5% * Me )that of theearth,
and its radius (Rv = 94.9% Re ) that of the earth.

a)The acceleration due to gravity on the surface ofVenus from these data

a = G * Mv / Rv 2
=G * ( 81.5 % Me) / [94.9% Re 2 ]
= [81.5 % / (94.9%) 2 ]* [G{Me}/( Re 2)]
= 0.90495 *g
= 8.87 m /s2
Here g = [G{Me}/ ( Re2)] = 9.81 m /s 2

b) The weight of a (m = 5.00 kg) rock on the surface ofVenus
W = mg
= 44.3 N
answered by: neel